OpenStudy (anonymous):

What is the value of x? a) 5 b) 2.5 c) 7.5 d) 10

OpenStudy (anonymous):

@undeadknight26 and @Mr.Man420

OpenStudy (anonymous):

not sure here sorry.

OpenStudy (campbell_st):

the triangles are similar, so the ratio of corresponding sides are equal  use the ratio of large triangle to small triangle \[\frac{2x + 5}{x} = \frac{2x -1}{x - 2} \] so now cross multiply, collect like terms and solve for x. hope it makes sense

OpenStudy (anonymous):

ok hold one give me time

OpenStudy (anonymous):

5

OpenStudy (campbell_st):

that's correct

OpenStudy (anonymous):

thanks!!!